### v3

# Steering Away from Harm: An Adaptive Approach to Defending Vision Language Model Against Jailbreaks

Abstract
Vision Language Models (VLMs) can produce unintended and harmful content when exposed to adversarial attacks, particularly because their vision capabilities create new vulnerabilities. Existing defenses, such as input preprocessing, adversarial training, and response evaluation-based methods, are often impractical for real-world deployment due to their high costs. To address this challenge, we propose ASTRA, an efficient and effective defense by adaptively steering models away from adversarial feature directions to resist VLM attacks. Our key procedures involve finding transferable steering vectors representing the direction of harmful response and applying adaptive activation steering to remove these directions at inference time. To create effective steering vectors, we randomly ablate the visual tokens from the adversarial images and identify those most strongly associated with jailbreaks. These tokens are then used to construct steering vectors. During inference, we perform the adaptive steering method that involves the projection between the steering vectors and calibrated activation, resulting in little performance drops on benign inputs while strongly avoiding harmful outputs under adversarial inputs. Extensive experiments across multiple models and baselines demonstrate our state-of-the-art performance and high efficiency in mitigating jailbreak risks. Additionally, ASTRA exhibits good transferability, defending against unseen attacks (i.e., structured-based attack, perturbation-based attack with project gradient descent variants, and text-only attack). Our code is available at https://github.com/ASTRAL-Group/ASTRA.
1 Introduction
Vision Language Models (VLMs) [28, 12, 61, 8] have attracted significant attention from both the industry and academia for their remarkable vision-language cognition capabilities [39]. Despite widespread applications, VLMs still face safety challenges due to limitations inherent in their underlying language models. Moreover, integrating visual inputs can open up a new surface for adversarial attacks. These safety issues regarding VLM have led to a lot of research on jailbreak attacks and defense strategies [47, 62, 53, 17].
Jailbreak attacks in VLMs aim to induce models to generate harmful responses by using jailbreaking image-text pairs [24, 23, 22, 51, 48, 41, 26]. These jailbreak attacks can be categorized into two types: (i) perturbation-based attacks, which create adversarial images that prompt generation of the harmful response from VLMs [41, 46, 2, 38], (ii) structured-based attacks, which embeds the malicious queries into images via typography to bypass the safety alignment of VLMs [17, 30]. Countermeasures for both attacks have been explored extensively: the input preprocessing-based method [37] or adversarial training [25] have proven effective for perturbation-based attacks. However, these defenses suffer as they require intensive computations to purify the image or fine-tune the model. Response evaluation-based  [18, 53, 58] defenses have been proposed for structured-based attacks, but they all require running model inference multiple times to potentially identify harmful outputs, which dramatically increases the cost of real-world deployment.
In this work, we argue that an efficient defense framework should not require significant computational resources during training or generating responses multiple times during inference. Drawing inspiration from recent advancements in activation steering in Large Language Model (LLM) [20, 43, 52, 4], we propose ASTRA, an efficient and effective defense by adaptively steering models away from adversarial feature directions via image attribution activations to resist VLM attacks. We find that simply borrowing the method from steering LLM for safeguarding VLM is not empirically workable due to the mismatch between the steering vectors obtained from textual and visual data, which necessitates our image attribution approach.
Specifically, ASTRA consists of two steps: constructing steering vectors via image attribution, and adaptive activation steering at inference time. We seek to construct steering vectors representing the direction of harmful responses. This can be done by constructing a set of adversarial images (e.g., using projected gradient descent (PGD) [34] algorithm) and then identifying visual tokens in each adversarial image most likely to trigger the jailbreak. To attribute such visual tokens, we fit a linear surrogate model using Lasso and estimate the impact of the inclusion/exclusion of each visual token on the probability of jailbreaks. The top-$k$ impactful visual tokens are then used to construct the steering vectors. This surrogate can be quickly estimated with only a few inference passes, making the process of building defense computationally friendly. During inference, we propose adaptive steering to manipulate the model’s activation through an activation transformation step. The steering coefficient is determined by the projection between the calibrated activation and steering vector, making the steering have little effect on benign input and a strong effect on adversarial input. This process is also efficient since it only requires generating a single response.
Extensive experiments demonstrate that ASTRA effectively mitigates perturbation-based attacks while preserving model utility across standard VLM benchmarks. The main contributions of this work are as follows:
  • $\circ$
    We introduce ASTRA, a defense that adaptively steers models away from adversarial feature directions via image attribution activations to resist VLM attacks. ASTRA is also highly efficient, which only needs several times of inference passes to build the defense, and does not affect inference time deploying the defense.
  • $\circ$
    We propose an adaptive steering approach by considering the projection between the steering vectors and calibrated activations, resulting in little performance drops on benign inputs while strongly avoiding harmful outputs under adversarial inputs.
  • $\circ$
    ASTRA achieves a substantial improvement in defending against perturbation-based attacks. Compared to state-of-the-art methods JailGuard [58], with a Toxicity Score of 12.12% and an Attack Success Rate of 17.84% lower, and 9x faster in MiniGPT-4. ASTRA is also transferable to some unseen attacks (i.e., structure-based attack, perturbation-based attack with PGD variants, and text-only attack), and still be effective against adaptive attacks.
2 Related Work
Jailbreak Attacks on VLM. Jailbreak attacks aim to alter the prompt to trick the model into answering forbidden questions. Apart from the LLM-based textual jailbreak strategies [31, 64, 19, 56, 60], additional visual inputs expose a new attack surface to VLM attacks. There are two main types of attacks: perturbation-based attacks and structured-based attacks [53]. Perturbation-based attacks create adversarial images to bypass the safeguard of VLMs [6, 41, 2, 46, 55, 59]. Structued-based attacks convert the harmful content into images through typography or text-to-image tool (e.g., Stable Diffusion [44]) to induce harmful responses from the model [29, 17, 30, 33, 27] We study our defense on both types of attacks.
Defenses on VLM. Researchers have explored two directions for defense: training-time alignment and inference-time alignment. Training-time alignment safeguards VLMs through supervised fine-tuning (SFT) [62, 9, 26] or training a harm detector to identify the harmful response [40], all requiring considerable high-quality annotation and sufficient computation resources to train. Inference-time alignment is relatively more resource-friendly. Some strategies design alignment prompts to defend against attacks [17, 54]. Others build a response evaluation pipeline to assess the harmfulness of VLM responses, often followed by iterative refinement to ensure safe outputs [18, 58]. Another way is to disturb input queries and analyze response consistency to identify potential jailbreak attempts [58]. However, these methods still introduce a non-trivial cost to inference time due to the need for generating the response multiple times.
Activation Engineering of LLM. The activation space of many language models appears to contain interpretable directions, which play a crucial role during inference [5, 36]. The basic idea of activation engineering is to identify a direction (i.e., steering vector) in activation space associated with certain semantics and then shift activations in that direction during inference. Turner et al. [50] locates the direction by taking the difference in intermediate activations of a pair of prompts at a particular layer and token position in a transformer model. Rimsky et al. [43] construct a dataset of contrast pairs rather than using a single pair to get the steering vector. Wang et al. [52] locate the “safety” steering vectors from a well-aligned language model. Ball et al. [4] investigate whether different types of jailbreak templates employ distinct mechanisms to trigger unsafe regions in the model’s representation space. Some other methods try to learn high-level concepts in the representation space and use them to control the output [63, 65, 20]. However, most previous works focus on utilizing textual prompts to construct steering vectors, which might not be empirically workable for steering VLM in some cases due to the gap between visual and textual domains.
3 Methodology
Figure 1: Illustration of our framework ASTRA. Our key procedures involve finding transferable steering vectors representing the direction of harmful response and applying adaptive activation steering to remove these directions at inference time. To create effective steering vectors, we randomly ablate the visual tokens from the adversarial images and identify those most strongly associated with jailbreaks. These tokens are then used to construct steering vectors. During inference, we perform an adaptive steering method that involves the projection between the steering vectors and calibrated activation, resulting in little influence on benign inputs and a strong impact on adversarial inputs. The solid and dotted lines denote the activations $h^{l}$ and calibrated activations $h^{l}-h_{0}^{l}$ respectively. The blue refers to the calibration activation $h_{0}^{l}$. The color red denotes the case of adversarial inputs.
In this work, we propose ASTRA, an efficient and effective defense by adaptively steering (Section 3.2) models away from adversarial directions via image attribution activations (Section  3.1) to resist VLM attacks.
Notation. Let $\mathcal{P}_{\text{VLM}}$ be an autoregressive vision language model, which defines a probability distribution over a sequence of preceding tokens from a vocabulary $\mathcal{V}$. Specifically, we consider a VLM which takes a sequence of $n$ textual tokens $\mathbf{x}_{t}=\{x_{t_{1}},...,x_{t_{n}}\}$ and $m$ visual tokens $\mathbf{x}_{v}=\{x_{v_{1}},...,x_{v_{m}}\}$ to generate responses $\mathbf{r}=\{r_{1},...,r_{o}\}$. We generate the $i\text{th}$ token $r_{i}$ of the response as follows:
$r_{i}\sim\mathcal{P}_{\text{VLM}}(\cdot\mid x_{v_{1}},\dots,x_{v_{m}},x_{t_{1}},\dots,x_{t_{n}},r_{1},\dots,r_{i-1})$
Algorithm 1 Pipeline of constructing steering vectors
Input: VLM $\mathcal{P}_{\text{VLM}}$, a set $\mathcal{D}$ of adversarial visual tokens $\mathbf{x}_{v}$, harmful instruction tokens $\mathbf{x}_{t}$, number of ablations $N$, template tokens $\mathbf{x}_{\text{template}}$, $\textbf{a}^{l}(\cdot)$ is the activation of layer $l$ in the VLM
Initialize $i\leftarrow 0$, $n\leftarrow 0$, specify $\mathbf{r}$ as tokens of “Sure, …”
while $i<|\mathcal{D}|$ do
    $n\leftarrow 0$
    while $n<N$ do
         Compute: $f(g_{n})=\text{log}\mathcal{P}_{\text{VLM}}(\mathbf{r}|\text{Ablate}(\mathbf{x}_{v},g_{n}),\mathbf{x}_{t})$
         $n\leftarrow n+1$
    Fit a linear surrogate model $\hat{f}$ using Lasso based on the pairs of $\{(g_{1},f(g_{1})),\dots,(g_{N},f(g_{N}))\}$
    Mask the visual tokens with the Top-$k$ weights in the $\hat{f}$ and get Mask($\mathbf{x}_{v}$)
    Construct the steering vector $v^{l}_{i}=\mathbf{a}^{l}(\mathbf{x}_{v},\mathbf{x}_{\text{template}})-\mathbf{a}^{l}(\text{Mask}(\mathbf{x}_{v}),\mathbf{x}_{\text{template}})$
    $i\leftarrow i+1$
Average across the set $v^{l}=\sum_{i=0}^{|D|}v^{l}_{i}$
Output: steering vector $v^{l}$
Figure 2: Illustration of steering. The colors red and green denote the activations for adversarial and benign inputs. The colors blue and brown denote the calibration activations $h^{l}_{0}$ and steering vectors $v^{l}$.
3.1 Constructing Steering Vectors
Not all visual tokens from the adversarial images contribute to the jailbreak equally. We seek to locate certain visual tokens that have a higher chance of inducing jailbreaking via image attribution. In this way, we can isolate the representation most associated with jailbreak-related information in these tokens.
Adversarial Image Attribution. Image attribution aims to find the input visual tokens that are more likely to trigger the specified responses. In our case, we seek to locate visual tokens from adversarial images generated by the PGD attack with a higher chance of inducing the jailbreak.
We conduct random ablation of certain tokens and compute the impact of exclusion/inclusion on inducing the jailbreak. We define visual token ablation as the process of masking specific tokens in a visual input. Let $\text{Ablate}(\mathbf{x}_{v},g)$ represent ablated visual tokens $\mathbf{x}_{v}$, where $g\sim\{0,1\}^{m}$ is an ablation vector that designates which tokens to mask (zeros in $g$ indicate masked tokens). Given an ablation vector $g$, the image attribution is expected to quantify the impact on the log probability of generating specified responses $\mathbf{r}$,
$f(g):=\text{log}\mathcal{P}_{\text{VLM}}(\mathbf{r}|\text{Ablate}(\mathbf{x}_{v},g),\mathbf{x}_{t}),$
changes as a function of $g$, where $\mathbf{x}_{t}$ are textual tokens of harmful instructions, $\mathbf{r}$ as tokens of “Sure, $\dots$” to denote the case of jailbreaking, and $\mathcal{P}_{\text{VLM}}(\mathbf{r}\mid\text{Ablate}(\mathbf{x}_{v},g),\mathbf{x}_{t})$ as the product of the probability of generating specified response $\mathbf{r}$ given the $\text{Ablate}(\mathbf{x}_{v},g),\mathbf{x}_{t}$.
Following prior work in machine learning explanation [42, 10], we fit a linear surrogate model $\hat{f}$ to analyze the influence of masking subsets of visual tokens on the likelihood of jailbreaks and select the visual tokens that are highly relevant for triggering the jailbreaking responses. Specifically, we (1) sample a dataset of ablation vectors $g_{1},\dots,g_{n}$ and compute $f(g_{i})$ for each $g_{i}$ by multiple times of ablations and forwards, (2) train the surrogate model $\hat{f}:\{0,1\}^{m}\rightarrow\mathbb{R}$ using Lasso to approximates $f$ based on the pairs $(g_{i},f(g_{i}))$, and (3) attribute the behavior of the surrogate model $\hat{f}$ to individual visual tokens. Finally, we can get a surrogate model $\hat{f}$ with its weights that can be interpreted as the attribution scores for triggering the jailbreak. The higher the score, the more relevant the token results in jailbreak.
Harmful Feature Extraction. With attribution scores for each token, we extract the representation of those tokens strongly correlated with jailbreak. Additionally, we hope our steering vectors generalize rather than overfitting to specific instructions and enjoy good transferability to a wider range of jailbreaks. Thus, we utilize visual tokens with Top-$k$ attribution scores from the surrogate model $\hat{f}$ paired with the empty user query to construct the steering vectors.
Given a set $\mathcal{D}$ of ($\mathbf{x}_{v}$, $\text{Mask}(\mathbf{x}_{v})$) and textual tokens $\mathbf{x}_{\text{template}}$ of chat template with an empty user query, where $\mathbf{x}_{v}$ is the input visual tokens, and $\text{Mask}(\mathbf{x}_{v})$ is input visual tokens masked with Top-$k$ attributed tokens, we calculate the mean difference vector for a layer $l$ as:
$v^{l}=\frac{1}{|\mathcal{D}|}\sum_{\mathbf{x}_{v},\text{Mask}(\mathbf{x}_{v})\in\mathcal{D}}$ $\left[\mathbf{a}^{l}\left(\mathbf{x}_{v},\mathbf{x}_{\text{template}}\right)\right.$
$\left.-\mathbf{a}^{l}\left(\text{Mask}(\mathbf{x}_{v}),\mathbf{x}_{\text{template}}\right)\right]$
where $\mathbf{a}^{l}$ captures the activations at the last token in layer $l$. The difference between these pairs isolates the representation most associated with jailbreak-related information in visual tokens with Top-$k$ attribution scores.
3.2 Adaptive Activation Steering
[…]
Benign image 30.65 30.65 30.65 30.65 24.55 24.55 24.55 24.55
Adversarial image $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained
VLM defenses
w/o defense 39.73 48.52 54.70 52.12 44.55 47.27 49.09 53.64
Self-reminder [54] 38.97 48.71 45.15 50.12 35.45 36.36 41.82 43.64
JailGuard [58] 16.51 18.93 20.93 21.23 30.00 32.73 27.27 28.18
ECSO [18] 34.59 32.42 38.54 42.86 40.91 42.73 29.09 37.27
LLM Steering
Refusal Pairs [43] 25.76 30.28 31.99 35.71 20.00 22.73 17.27 16.36
Jailbreak Templates [4] 19.73 25.03 30.10 22.78 33.64 38.15 38.18 42.73
ASTRA (Ours) 11.30 8.84 4.51 4.48 9.09 13.18 15.46 9.09
[…]
Benign image 38.52 38.52 38.52 38.52 0.00 0.00 0.00 0.00
Adversarial image $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained
VLM defenses
w/o defense 50.50 51.62 55.59 53.43 67.27 70.46 71.82 76.36
Self-reminder [54] 30.47 27.53 32.84 29.09 50.00 47.27 40.00 58.18
JailGuard [58] 29.37 24.68 28.74 27.76 19.09 20.00 21.82 15.45
ECSO [18] 50.09 50.68 56.08 51.57 30.00 27.27 31.82 32.73
LLM Steering
Refusal Pairs [43] 46.14 46.83 46.83 40.53 29.09 31.82 21.82 52.73
Jailbreak Templates [4] 66.74 63.35 67.15 68.29 68.18 68.18 65.45 74.55
ASTRA (Ours) 15.52 5.45 2.39 0.07 6.06 5.00 18.18 15.45
The key idea of activation steering is using steering vectors to shift a language model’s output distribution toward a specified behavior during inference. After constructing steering vectors with harmful semantics, we strive to remove these components by steering LLM’s activations.
Unfortunately, simply applying a fixed scaling coefficient to the steering vector for modifying the language model’s output [43, 4, 52, 50] is not workable as a defense due to dramatic utility performance degradation in benign cases [1]. The main problem is that the linear steering used in prior work unconditionally alters the activation no matter whether the input leads to harmful outputs or not (Fig. 2(a)):
$h^{l}=h^{l}-\alpha\cdot\frac{v^{l}}{\|v^{l}\|}$
where $h^{l}$ is the activation of the last token at the layer $l$, and $\alpha$ is a scaling coefficient. To address this challenge, we propose adaptive steering based on conditional projection:
$h^{l}=h^{l}-\alpha\cdot\text{max}(\frac{(h^{l})^{\top}v^{l}}{\|h^{l}\|\|v^{l}\|},0)\cdot\frac{v^{l}}{\|v^{l}\|}$
When $h^{l}$ does not contain any positive component of the steering vector (harmful direction), the $\max$ term is 0, leaving activations unchanged. This minimized the negative impact on the benign performance.
Since the angle between $h^{l}$ and $v^{l}$ matters for adaptive projection, we must ensure that it can effectively distinguish harmful and benign activations at layer $l$. However, we notice that the activations for different inputs may cluster around a point distant from the origin. As a result, the angles among these vectors may all become similar (Fig. 2(b)). To address this, we propose a activation calibration step before steering. We use the calibration activation $h_{0}^{l}$, which can be seen as the center of the activation for many different inputs, to calibrate the projection term in our adaptive steering:
$h^{l}=h^{l}-\alpha\cdot\text{max}(\frac{(h^{l}-h^{l}_{0})^{\top}v^{l}}{\|h^{l}-h^{l}_{0}\|\|v^{l}\|}\cdot\|h^{l}\|,0)\cdot\frac{v^{l}}{\|v^{l}\|}$
$h_{0}^{l}$ is the calibration activation at the layer $l$, $h^{l}-h_{0}^{l}$ is the calibrated activation. We do not calibrate $v^{l}$ here since the mean component has been canceled out when subtracting the two token activations. To obtain the calibration activation $h_{0}^{l}$, we collect image-text queries from a large number of test data and compute the average of the generated token features at the layer $l$ to get $h_{0}^{l}$.
We show the full process of our adaptive steering approach in Fig. 2 (c1) - (c3). It can help reduce malicious outputs in adversarial scenarios while preserving performance in benign cases. During inference, we apply steering only to the activations of newly generated tokens, leaving the activations of input tokens unaltered.
4 Experiments
In this section, we conduct experiments to address the following research questions:
  • $\bullet$
    RQ1: How does ASTRA perform in adversarial scenarios compared to VLM defense baselines and LLM steering methods? Is our defense transferable to a different distribution of inputs and different types of attacks?
  • $\bullet$
    RQ2: How does ASTRA perform in benign cases? Can we reduce model harmfulness without hurting utility?
  • $\bullet$
    RQ3: What are the impacts of design choices in ASTRA? Are all components (e.g., image attribution, activation calibration) necessary for best performance?
Table 3: The performance against adaptive attacks. The adversary has complete knowledge of the model, our steering vectors and adaptive steering defense mechanism. Under this strong (often unrealistic) attack setting, ASTRA still noticeably outperform undefended models.
[…]
$\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained
MiniGPT-4
Attack on undefended VLM 39.73 48.52 54.70 52.12 44.55 47.27 49.09 53.64
Adaptive Attack on defended VLM 15.47 19.23 20.50 17.04 13.64 13.64 24.55 22.73
Qwen2-VL
Attack on undefended VLM 50.50 51.62 55.59 53.43 67.27 70.46 71.82 76.32
Adaptive Attack on defended VLM 24.56 24.21 9.27 11.60 58.16 60.00 59.09 69.09
LLaVA-v1.5
Attack on undefended VLM 83.70 84.40 85.54 85.44 51.82 56.36 55.45 53.64
Adaptive Attack on defended VLM 60.24 63.59 68.87 67.86 30.00 34.55 32.73 32.73
Table 4: Inference Time per token (ms). “Single inference” indicates whether the method requires generating responses multiple times during evaluation. We report inference time per token since the total inference time may vary depending on the length of the generated tokens.
Single Inference Toxicity (Perturbation-based Attack)
MiniGPT-4 LLaVA-v1.5 Qwen2-VL
w/o defense 173.19 40.68 27.43
Self-reminder [54] 173.36 41.09 27.94
JailGuard [58] 1557.98 366.02 245.42
ECSO [18] 457.55 116.44 70.22
ASTRA (Ours) 173.77 40.69 27.98
Figure 3: Transferability in ID scenarios. Avg. denotes the average of steering vectors derived from the adversarial images with $\epsilon$ values in $\frac{16}{255},\frac{32}{255},\frac{64}{255}$, unconstrained}. Additional results for LLaVA-v1.5 can be found in Appendix, Fig. 7.
Table 5: Transferability in OOD scenarios. We evaluate the transferability of steering vectors derived from the Jailbreak adversarial images with $\epsilon=\frac{16}{255}$ and choose the same $\alpha$ tuned on the Jailbreak validation set. The transferability is evaluated across multiple unseen attack categories: structured-based attack from MM-SafetyBench [29], perturbation-based attack with various PGD variants, and text-only attack. We use the classifier from HarmBench [35] to compute the ASR.
Structured-based Attack Perturbation-based Attack Text-only Attack
SD SD_TYPO TYPO PGD [34] Auto-PGD [11] MI-FGSM [14] GCG [64]
$\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=16/255$ $\epsilon=32/255$
MiniGPT-4
w/o defense 13.75 43.25 43.75 70.91 78.18 74.55 76.36 78.18 79.09 58.18
ASTRA (Ours) 3.75 8.75 11.25 5.45 12.73 5.45 10.91 16.37 13.64 9.09
Qwen2-VL
w/o defense 20.00 61.25 38.75 74.55 80.00 76.37 77.57 80.00 78.18 81.82
ASTRA (Ours) 11.25 40.00 33.75 21.82 14.55 15.76 15.76 18.18 18.18 30.91
LLaVA-v1.5
w/o defense 18.75 55.00 22.50 69.09 74.55 80.60 90.30 87.28 89.09 92.73
ASTRA (Ours) 8.75 25.00 6.25 1.82 1.82 1.21 0.61 0.00 0.00 14.55
4.1 Experimental Setup
Steering Vector Construction. We sample benign images with different classes from ImageNet [13] and apply the PGD attack [34] to generate 16 adversarial images for steering vectors construction. The perturbation radius $\epsilon$ is set to $\{\frac{16}{255},\frac{32}{255},\frac{64}{255},\text{unconstrained}\}$. Details on the PGD attack configuration can be found in Appendix 8.1.
Evaluation Datasets. We choose Toxicity and Jailbreak setups using the perturbation-based attack. We sample 55 benign images from ImageNet [13] and apply the PGD attack [34] to generate 25 and 30 adversarial images for visual validation, and test sets respectively. The perturbation radius $\epsilon$ is set to $\{\frac{16}{255},\frac{32}{255},\frac{64}{255},\text{unconstrained}\}$. For textual prompts, we choose 50 and 100 queries from RealToxicityPrompt [16] to construct the validation and test set for Toxicity setup. We choose 110 and 110 queries from both Advbench [64] and Anthropic-HHH [15] to construct the validation and test set for Jailbreak setup. All text prompts are different from the instruction-response pairs used for steering vector construction. During the evaluation, we pair each textual prompt with a random adversarial image. Detailed data information can be found in Fig. 4.
For the evaluation of utility performance in benign scenarios, we employ two established benchmarks, MM-Vet [57] and MM-Bench [32]. Additionally, we include safe instructions from XSTest [45] to assess the overrefusal case. Full Details of dataset statistics can be found in Appendix 8.1.
Evaluation Metrics. For the toxicity setup, we follow Qi et al. [41] and use the Detoxify classifier [21] to calculate the toxicity score. We report the average scores of Toxicity attribute across the test set. The scores range from 0 (least toxic) to 1 (most toxic). For the jailbreak setup, we choose the classifier from HarmBench [35] to compute the attack success rate (ASR).
Baselines. We compare ASTRA with three VLM defense baselines and two LLM steering approaches. For the VLM defenses, self-reminder [54] is a system prompt based defense, JailGuard [58] perturbs the input images several times and computes the divergence between responses, and ECSO [18] adaptively transforms unsafe images into texts to activate the intrinsic safety mechanism of pre-aligned LLM in VLMs. For the LLM steering, we follow Rimsky et al. [43] and Ball et al. [4] to construct steering vectors with the semantics of refusal and textual jailbreak templates.
Models & Implementations details. We conduct all the experiments on three popular open-sourced VLMs, including Qwen2-VL-7B [3], MiniGPT-4-13B [61], and LLaVA-v1.5-13B [28]. We set the number of ablations $N$ as 96, $k$ as 15. For the selection of $\alpha$, refer to Appendix 8.6. The steering layer $l$ is 20 for 13B models and 14 for 7B models. The chat configurations use a temperature of 0.2 and $p=0.9$ for LLaVA-v1.5 and Qwen2-VL, and a temperature of 1 and $p=0.9$ for MiniGPT-4.
4.2 Defense Performance Comparision (RQ1)
Table 12, and 8 (in appendix) report the performance of our defense in the perturbation-based attack across Toxicity and Jailbreak setup. Bold denotes the best defense performance (represented by Toxicity Score or ASR).
Comparison with Existing VLM Defenses. As shown in Table 128, most VLM defenses struggle to consistently safeguard the model against perturbation-based attacks with different $\epsilon$. While most existing VLM defenses are based on pre- or post-processing model inputs or outputs, our adaptive steering approach effectively steers the internal model activations away from harmful contents, achieving state-of-the-art performance across almost all cases.
Additionally, we report the average inference time per token for each VLM defense baseline in Table 4. We emphasize two key benefits that lead to high efficiency: (1) ASTRA does not need to re-train or fine-tune the model, and the process of constructing steering vectors (Section 3.1) is cheap and straightforward. In contrast, input preprocessing-based method [37] needs to denoise each input image using the Diffusion model and adversarial training [25] needs to update the entire model, both are quite costly compared to our approach. (2) ASTRA does not affect inference time when deploying the defense - the steering step in Section 3.2 has almost negligible cost. As shown in Table 4, ASTRA are faster than those methods requiring multiple inference passes (e.g., JailGuard [58] and ECSO [18]). While JailGuard [58] can defend against perturbation-based attacks effectively, it requires generating nine responses to deploy the defense and can be highly costly. While self-reminder [54] does not impact inference time, it fails to protect VLMs against perturbation-based attacks in most cases.
Overall, these empirical results validate both the effectiveness and efficiency of our framework in defending against VLM perturbation-based attacks.
Table 6: Utility performance in benign and adversarial scenarios. “Direct” denotes the performance of original VLMs. Bold=better.
[…]
MM-Vet [57] MMBench [32] XSTest [45] Toxicity (Perturbation-based) Jailbreak (Perturbation-based) Jailbreak (Structured-based)
Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA Direct ASTRA
MiniGPT-4 19.40 20.62 35.90 35.82 87.60 87.60 51.42 10.14 3.95 5.82 2.62 4.29
LLaVA-v1.5 32.62 30.55 72.94 73.23 98.00 98.80 63.68 59.28 3.68 8.59 3.82 4.61
Qwen2-VL 49.13 48.66 78.00 78.69 73.60 74.00 140.44 40.14 6.80 8.86 30.00 30.92
Table 7: Ablation study of adaptive steering on Qwen2-VL. “Random Noise” means steering with Gaussian noise, “Entire Img” refers to steering with the entire image activation, “Img Attr” represents steering using the image attribution activation, and “Calibration Activation” indicates whether the calibration activation is incorporated into the projection term.
[…]
Steering Vector Calibration Activation $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained
Random Noise 44.10 53.80 61.09 55.10 64.55 67.27 69.09 76.36
Entire Img 42.60 44.40 49.61 29.53 60.91 44.55 63.64 75.45
Img Attr 40.49 41.80 33.90 10.50 50.00 24.55 51.82 72.73
Entire Img 37.70 35.28 21.59 5.24 46.82 47.28 42.73 22.73
Img Attr (Ours) 15.52 5.45 2.39 0.07 6.06 5.00 18.18 15.45
Comparison with LLM Steering. Our results in Table 128 indicate that directly adapting steering techniques from LLMs to VLM defenses is ineffective. While steering vectors infused with refusal semantics can shift output distribution toward refusal and lower harmful response rates, this approach has a critical drawback: it indiscriminately increases refusal rates across all inputs, which diminishes model utility [1]. Furthermore, our experiments reveal that steering with textual jailbreak templates is insufficient to counteract perturbation-based attacks on images, suggesting that textual and visual jailbreaks exploit different mechanisms to circumvent VLM safeguards. These findings emphasize the importance of developing VLM defenses that operate at the visual representation level.
Adaptive Attack. Adaptive attack [49] is a critical evaluation procedure for assessing defense effectiveness when the defense mechanism is known to the attacker. In this setup, we assume the attacker can access the model parameters, steering vector $v^{l}$, the calibration activation $h^{l}_{0}$, and steering coefficient $\alpha$, and employs the PGD attack to generate 30 adversarial images specifically targeting the defended model. As shown in Table 3, ASTRA continues to provide robust protection for the VLM in most cases. These findings emphasize the potential of our method as a practical and resilient defense mechanism in real-world applications.
Transferability. In real-world scenario, unknown types of adversarial attacks highlight the need for a robust and transferable defense framework. To evaluate transferability of ASTRA, we construct two test cases: in-distribution (ID) and out-of-distribution (OOD).
In ID scenario, adversarial images used for steering vector construction and test evaluations are drawn from same classes in ImageNet [13], ensuring similar image distributions. We assess whether steering vectors derived from adversarial images with a specific $\epsilon$ value can defend against adversarial images with varying $\epsilon$ levels. As illustrated in Fig. 3 and 7, the results demonstrate the effectiveness of our steering vectors defending against adversarial attacks with different $\epsilon$ values. We also report the Avg. performance, in which we take the mean of steering vectors derived from adversarial images with $\epsilon$ values in $\{\frac{16}{255},\frac{32}{255},\frac{64}{255},\text{unconstrained}\}$. Despite that the defense with $\epsilon=\text{unconstrained}$ does not work quite well against perturbation-based attacks with $\epsilon=\{\frac{16}{255},\frac{32}{255},\frac{64}{255}\}$, remaining defense validate the transferability of ASTRA across PGD attacks with different intensities.
In OOD scenario, we test whether steering vectors derived from the Jailbreak adversarial images with $\epsilon=\frac{16}{255}$ can generalize to different types of attacks. Specially, we evaluate the defense transability on structured-based attack from MM-SafetyBench [29], perturbation-based attack with several PGD variants, and text-only attack. Please refer Appendix 8.1 for details of structured-based attack. For the perturbation-based attack, we collect 12 images with distributions differing from images used for steering vector construction (e.g., stripes, sketch, painting, etc). We use 55 instruction-response pairs from JailbreakBench [7] to conduct perturbation-based attacks with PGD variants (i.e, PGD, MI-FGSM [14], and Auto-PGD [11]) and text-only attack (i.e., GCG [64]). We use the same 55 instructions from JailbreakBench [7] to evaluate performance.
Results in Table 5 confirm the defense transferability across different unseen attacks, indicating great potential for real-world deployment. This impressive OOD transferability may arise from the steering vectors encapsulating a common harmful feature direction that persists regardless of how the harmful behavior is triggered. Although models can be jailbroken by different types of attacks, eventually, there exists a certain direction in the feature space that represents the harmfulness. By accurately steering away from this direction, we can effectively safeguard models against diverse types of jailbreaks.
4.3 General Utility (RQ2)
In Section 4.2, our framework demonstrates its effectiveness in defending against VLM jailbreaks. Furthermore, we need to ensure that our defended model retains utility performance in benign scenarios and generates valid responses in adversarial scenarios.
Utility Performance. We calculate utility scores in MM-Vet [57], MMBench [32], and safe instructions from XSTest [45] for benign scenario evaluation and perplexity for adversarial scenario evaluation. See Appendix 8.1 for detailed descriptions of utility scores. As shown in Table 6, our defended models demonstrate considerable utility performance in benign scenarios compared to those without defenses. These comparisons demonstrate that our defense results in little performance drops on benign inputs. We owe these results to our adaptive steering approach, which mitigates utility degradation by computing the projection between the language model’s calibrated activation and steering vectors, thereby avoiding the drawbacks of a fixed steering coefficient. In adversarial contexts, the perplexities of ASTRA are still within a reasonable range, indicating that our defended models consistently provide valid, non-harmful responses to harmful instructions. Additional cases are provided in Appendix 8.4.
4.4 Ablation Study (RQ3)
Adaptive Steering. We demonstrate the roles of calibration activation and image attribution in our adaptive steering operation using Qwen2-VL. As shown in table 7, both designs significantly influence defense performance. Specifically, after calibration activation, the projection term can more accurately reflect the spatial relationship between steering vectors and activations within the feature space, leading to a consistent defense effectiveness in both Toxicity and Jailbreak setups. Furthermore, we compare the performance of steering vectors derived from the image attribution activation versus those derived from the entire image activation. Steering vectors from the entire image are constructed by averaging $\mathbf{a}^{l}(\mathbf{x}_{v},\mathbf{x}_{\text{template}})-\mathbf{a}^{l}(\mathbf{x}^{\text{empty}}_{v},\mathbf{x}_{\text{template}})$ across the set of 16 adversarial images for vector construction, where $\mathbf{x}_{v}$ is the adversarial image, $\mathbf{x}_{\text{template}}$ is the chat template, and $\mathbf{x}^{\text{empty}}_{v}$ is an empty image. The results demonstrate the importance of our image attribution procedure. By narrowing down to certain visual tokens strongly associated with the jailbreak behavior, our image attribution better isolates jailbreak-related information. We also conducted experiments using random noise vectors to assess the potential influence of noise on our framework. These results suggest that steering with image attribution activations offers superior performance compared to steering with entire image activations or random noise, providing a more targeted and effective defense mechanism.
Please refer to Appendix 8.5 for more ablation studies on steering coefficient $\alpha$, number of adversarial images used for steering vector construction, and steering layer selection.
5 Conclusion
In this paper, we propose ASTRA, an efficient and effective defense framework by adaptively steering models away from adversarial feature directions to resist VLM attacks. Our key procedures involve finding transferable steering vectors representing the direction of harmful response via image attribution and applying adaptive activation steering to remove these directions at inference time. Extensive experiments across multiple models and baselines demonstrate our state-of-the-art performance and high efficiency. We hope our work will inspire future research on applying more sophisticated steering for LLM/VLM safety.
6 Acknowledgment
This work was supported by NSF 2331967, 2229876 and 2055233. Huan Zhang is supported in part by the AI2050 program (Early Career Fellowship) at Schmidt Sciences.
Supplementary Material
Figure 4: Data used for steering vectors construction and test evaluation in the Jailbreak setup.
7 Input Prompts
We provide detailed visual and textual templates for prompting MiniGPT-4, Qwen2-VL, and LLaVA-v1.5, as shown in Fig. 5.
Figure 5: Prompt template.
Figure 6: Samples of adversarial images used for steering vector construction, their corresponding visual tokens with top-$k$ attributions scores, and adversarial images for OOD scenario evaluation.
8 Experiment Details and Extra Results
[…]
Benign image 75.00 75.00 75.00 75.00 13.64 13.64 13.64 13.64
Adversarial image $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained
VLM defenses
w/o defense 83.70 84.40 85.54 85.44 51.82 56.36 55.45 53.64
Self-reminder [54] 83.92 83.97 84.19 80.93 28.18 30.00 22.73 22.73
JailGuard [58] 77.60 77.77 75.76 73.76 23.64 21.82 30.00 17.27
ECSO [18] 73.77 73.14 71.32 66.81 24.55 21.82 14.55 20.00
LLM Steering
Refusal Pairs [43] 66.72 66.82 60.36 62.46 23.64 25.45 20.00 19.09
Jailbreak Templates [4] 52.61 50.21 55.48 54.90 23.64 17.27 20.00 29.09
Ours 36.02 34.76 43.13 25.10 4.55 10.91 13.64 12.43
Figure 7: Transferability in ID scenarios on LLaVA-v1.5. Avg. denotes the average of steering vectors derived from the adversarial images with $\epsilon$ values in $\frac{16}{255},\frac{32}{255},\frac{64}{255}$ unconstrained}. The numbers shown in the figures are toxicity scores (left) and attack success rates (right).
8.1 Dataset Statistics
Implementation details of PGD attack
We use the PGD attack to inject adversarial noise into each benign image. For the Jailbreak setup, we prepare 416 and 415 harmful instructions and corresponding affirmation responses from the AdvBench [64] and Anthropic-HHH [15] respectively to conduct the PGD attack. For the Toxicity setup, we choose 66 toxic queries from Qi et al. [41] as the optimization objective to conduct the PGD attack. We apply PGD for 2500 iterations with a step size of 1/255 on MiniGPT-4, and a step size of 1/1020 on Qwen2-VL and LLaVA-v1.5.
Harmful instructions used for steering vector construction
For the Jailbreak setup, we use the same 416 harmful instructions from AdvBench [64] as PGD attacks. For the Toxicity setup, we use 40 harmful instructions from Qi et al. [41]. These instructions explicitly ask for the generation of detrimental content across four distinct categories: identity attack, disinformation, violence/crime, and malicious behaviors toward the human race. We run three times of image attributions on each adversarial image paired with different harmful instructions when constructing the steering vectors.
Datasets of structured-based attacks
We choose MM-SafetyBench [32] to evaluate our defense performance on unseen attacks (i.e., structured-based attacks). In this dataset, images contain most of the malicious content, while the text queries are benign. The image can be from one of the following: (1) SD: generated by Stable Diffusion based on malicious keywords, (2) TYPO: embedding text in blank images, and (3) SD_TYPO: embedding text in the image generated by Stable Diffusion. We randomly sample 10 items from each scenario in 01-07 & 09 to construct the test set: 01-Illegal Activity, 02-HateSpeech, 03-Malware Generation, 04-Physical Harm, 05-Economic Harm, 06-Fraud, 07-Pornography, 09-Privacy Violence. In this case, we have 80 test items for each setup.
Utility Datasets
We choose two established datasets to evaluate the defended models’ utility performance: MM-Bench [32] and MM-Vet [57].
MM-Bench [32] evaluates twenty different vision language capabilities through single-choice questions. We randomly sample 100 items and 200 items from the dataset to construct our validation and test set, respectively. We compute the accuracy of all the questions as the utility score in this dataset.
MM-Vet [57] evaluates six core vision language capabilities of VLMs, including recognition, knowledge, optical character recognition, language generation, spatial awareness, and math. MM-Vet requires the VLM to answer the question in an open-ended manner, which is a more challenging task than single-choice questions. To evaluate the performance, MM-Vet [57] queries GPT-4 with few-shot evaluation prompts to obtain a utility score ranging from 0 to 1. We randomly sample 50 and 100 items from the dataset to construct our validation and test set, respectively. We average across the scores for each item as the utility score in this dataset.
XSTest [45] evaluates the exaggerated safety behavior in LLMs. To construct our test set, we use 250 safe prompts across ten prompt types that well-calibrated models should not refuse to comply with. We define the utility score as the proportion of safe prompts with which the model complies, measured via the string matching.
8.2 Calibration Activation
In Section 3.2, we introduce the calibration activation to calibrate the projection term in the adaptive steering. To construct the calibration activation, we collect 21 images from ImageNet [13] and pair them with the prompt “What is the image about?”. We use these pairs to query the VLM and store activations of generated tokens at the layer $l$. Then, we average these activations to get the calibration activation $h_{0}^{l}$.
8.3 Extra Results on LLaVA
We provide additional quantitative results on the defense performance comparison on LLaVA-v1.5 in Table. 8 and transferability in ID scenarios on LLaVA-v1.5 in Fig. 7. These empirical results also demonstrate the effectiveness of our defense framework and the transferability across PGD attacks with different intensities.
8.4 Qualitative Results
Qualitative results for Qwen2-VL, LLaVA-v1.5, and MiniGPT-4 under adversarial scenarios are shown in Fig. 1011, and 12 respectively, while results under benign scenarios are provided in Fig. 13.
8.5 Ablation Study
(a) Linear Steering on MiniGPT-4
(b) Adaptive Steering on MiniGPT-4
(c) Linear Steering on Qwen2-VL
(d) Adaptive Steering on Qwen2-VL
Figure 8: Ablation study of steering coefficient and linear/adaptive steering in the Toxicity setup ($\epsilon=\frac{16}{255}$). MM-Vet [57] is used to evaluate performance in benign scenarios.
(a) Toxicity on LLaVA-v1.5
(b) Jailbreak on LLaVA-v1.5
(c) Toxicity on MiniGPT-4
(d) Jailbreak on MiniGPT-4
Figure 9: Ablation study of (a-b) the number of adversarial images used for steering vector construction on LLaVA-v1.5. (c-d) steering layer selection on MiniGPT-4. ($\epsilon=\frac{16}{255}$)
Steering Coefficient
We investigate the effect of steering coefficient $\alpha$ in the Toxicity setup using MiniGPT-4 and Qwen2-VL, comparing linear steering with our adaptive steering approach in both adversarial and benign scenarios. For utility evaluation, we use the more challenging MM-Vet dataset [57]. In the linear steering approach, the steering vector is normalized and multiplied by a fixed $-\alpha$ to steer the activation. As shown in Fig. 8(a) and (c), although linear steering performs well in adversarial scenarios, it struggles to maintain a considerable utility performance in benign scenarios. This imbalance between the defense and utility significantly limits its practical capability. This trend is also consistent with the insight in [1], emphasizing the need for an adaptive steering approach. As illustrated in Fig. 8(b) and (d), adaptive steering achieves a balance between defense and utility. We owe this balance to the projection term in our steering operation. By considering the projection between calibrated activations and steering vectors, our approach can effectively defend against adversarial attacks while preserving general performance in benign cases.
Number of adversarial images used for steering vector construction
We examine how the number of adversarial images used for steering vector construction affects defense performance using LLaVA-v1.5. As shown in Fig. 9(a) and (b), increasing the number of adversarial images for steering vector construction leads to rapid convergence in defense performance, indicating that only a modest amount of adversarial image is required. This result highlights the precision of our steering vectors in capturing the pattern of adversarial attacks.
Steering Layer Selection
We vary the selected steering layer to assess whether our framework can generalize across different layers. For simplicity, this ablation study uses linear steering, as it avoids tuning $\alpha$ for each layer. We multiply the normalized steering vector with the coefficient $\alpha=-0.8\|h^{l}\|$ to denote “Negative Steering” and $\alpha=+0.8\|h^{l}\|$ to denote “Positive Steering”. As shown in Fig. 9(c) and (d), we can shift the output semantics by selecting the appropriate middle or final layers. The results also indicate that our framework correctly identifies the harmfulness direction, enabling semantic manipulation through simple adjustments to the steering coefficient.
Figure 10: Qualitative examples on Qwen2-VL without defense (red) and with defense (green) in perturbation-based Toxicity, perturbation-based Jailbreak, and structured-based Jailbreak setups.
Figure 11: Qualitative examples on LLaVA-v1.5 without defense (red) and with defense (green) in perturbation-based Toxicity, perturbation-based Jailbreak, and structured-based Jailbreak setups.
Figure 12: Qualitative examples on MiniGPT-4 without defense (red) and with defense (green) in perturbation-based Toxicity, perturbation-based Jailbreak, and structured-based Jailbreak setups.
Figure 13: Qualitative examples of Qwen2-VL, LLaVA-v1.5, and MiniGPT-4 on MM-Vet [57] dataset. The colors red and green denote the original VLM and defended VLM respectively.
8.6 Implementation Details
Hyperparameter selections
We select steering coefficient $\alpha$ on the validation set, ensuring a balance between the utility scores and defense performance, shown in Table 9.
Computation infrastructure
All of the experiments are performed on a server with 9 NVIDIA L40S 48GB GPUs and two-socket 32-core Intel(R) Xeon(R) Gold 6338 CPU. The operating system is Ubuntu 22.04.5 LTS. The CUDA version is 12.6, the Python version is 3.10.14, and the Torch version is 2.4.0.
Table 9: Hyperparameter selection of steering coefficient $\alpha$.
Toxicity (Perturbation-based Attack) $\epsilon=16/255$ $\epsilon=32/255$ $\epsilon=64/255$ unconstrained
MiniGPT-4 5 5 5 5
Qwen2-VL 7 7 7 7
LLaVA-v1.5 10 10 10 15
Jailbreak (Perturbation-based Attack) $\epsilon=16$ $\epsilon=32$ $\epsilon=64$ unconstrained
MiniGPT-4 7 7 7 7
Qwen2-VL 7 7 7 7
LLaVA-v1.5 10 10 10 15